# Driving Referring Video Object Segmentation with Vision-Language Pre-trained Models

### v1

@@ IV-C Ablation studies
TABLE VI: Experimental results of cross-using VLP models and applying CLIP to other RVOS frameworks. All models are trained on Ref-Youtube-VOS alone.
Algorithms Pre-trained Pre-trained Aligned $\mathcal{J}\&\mathcal{F}$ (%)
@@
ReferFormer [7] Video-Swin-T RoBERTa 56.0 55.8
ReferFormer [7]+CLIP CLIP ViT-B/16 CLIP BERT 51.8 (-4.2) 51.1 (-4.7)
SgMg [60] Video-Swin-T RoBERTa 58.9 56.7
SgMg [60]+CLIP CLIP ViT-B/16 CLIP BERT 52.7 (-6.2) 51.9 (-4.8)
IV-C1 Analyses on proposed components.
We analyze the proposed components through 10 variants, as shown in Table IV. The experiments begin with a baseline (Variant-1) consisting of a frozen VLP model, $N$ transformer decoder layers originally used for stage-3 VL relation modeling, and a segmentation head.
Analyses on temporal-aware VL prompt-tuning. We gradually introduce different prompts into the baseline to analyze their effect (Variant-2/3/4). The language and vision prompts enable Variant-2 to adapt pre-trained representations for pixel-level prediction, improving $\mathcal{J}\&\mathcal{F}$ by 3.9%. By introducing the temporal prompts (3,072 learnable parameters) and the PRTC module (no learnable parameters), Variant-3 improves $\mathcal{J}\&\mathcal{F}$ by 2.3%. It shows that a few learnable parameters effectively enhance the temporal modeling ability. The performance gap between Variant-3 and Variant-4 indicates that historical prompts benefit RVOS in the clip-by-clip inference paradigm.
@@
Analyses on spatial-temporal attention. We construct Variant-8/9/10 modeling the dense spatial-temporal relation with our CF-MSA, global MSA, and 3D SW/W-MSA [9], respectively. Compared with Variant-7 which only models the temporal clue by the vision encoder, all the attention mechanisms bring performance gains, demonstrating the necessity of explicit spatial-temporal relation modeling. Besides, CF-MSA achieves the largest performance gain of 2.2%, demonstrating its effectiveness.
IV-C2 Analyses on adaptation methods
@@
We conduct experiments with several popular adaptation methods on VLP-RVOS to analyze their effect. The involved methods include: 1) Frozen, in which the VLP model is frozen; 2) Partial-$m$, in which the last $m$ layers of the vision encoder are fine-tuned; 3) Full fine-tuning, in which the entire vision encoder is fine-tuned (note that the language encoder is kept frozen following [7, 27, 15]); 4) Adapter-tuning, in which additional MLP layers are introduced to tune pre-trained representations. We use UniAdapter [72], which has proven to be effective on several cross-modality tasks.
@@
We report the within-dataset (Ref-Youtube-VOS) and cross-dataset (Ref-DAVIS17) evaluation results in Table V. Although the frozen method obtains the worst performance on Ref-Youtube-VOS, it performs well on Ref-DAVIS17 as it retains the pre-trained knowledge. Compared with partial fine-tuning, full fine-tuning cannot obtain better performance and even harm generalization. UniAdapter [72] obtains mediocre performance. We speculate the reason is that it is designed for image-level VL understanding. Our prompt-tuning performs best on the two benchmarks, demonstrating its effectiveness and generalization.
IV-C3 Analyses on temporal modeling methods
@@
We conduct experiments with two popular methods, TeViT [70] and IFC [71], which are also designed for temporal modeling within vision transformers. TeViT [70] shifts several learnable tokens across frames for temporal modeling. IFC [71] introduces a trainable transformer encoder for temporal aggregation, introducing 66 M learnable parameters. We evaluate the two methods in VLP-RVOS and report the results in Table V. The comparisons demonstrate the superiority of our PRTC. Moreover, PRTC occupies a small proportion of the computational load during the visual encoding process. For instance, the total FLOPs per frame for visual encoding with VLMo-L amount to approximately 152.2G, whereas for PRTC, it is only 1.9G.
IV-C4 Effect of the prompt token number
@@
IV-C5 Effect of the aligned VL space and our transferring framework
@@
We delve deeper into analyses by breaking the aligned VL space and applying CLIP to other RVOS frameworks. We break the aligned VL space by cross-using the vision and language encoders of CLIP and VLMo, which results in significant performance drops, as shown in Table VI. We also integrate CLIP into ReferFormer and SgMg, where we follow [73] to obtain hierarchical features based on ViT. Table VI shows that applying CLIP to ReferFormer or SgMg leads to large performance drops. These results highlight that both the aligned VL space and our transferring framework are crucial for achieving state-of-the-art performance.
IV-C6 Temporal modeling across different time spans
As shown in Figure 9, our VLP-RVOS exhibits superior and more consistent performance with varying inference clip lengths, in contrast to ReferFormer which uses a video backbone. The comparisons demonstrate the strong temporal modeling ability of our method.
Figure 9: Experimental results with varying inference clip lengths of ReferFormer (Video-Swin-T) and VLP-RVOS (ViT-B/16 CLIP), which have similar computation complexity. These models are trained on Ref-Youtube-VOS alone.
@@ IV-D Qualitative results
Figure 11 shows the segmentation results in a video with cluttered scenes. Our VLP-RVOS accurately comprehends detailed descriptions and precisely segments the referred objects, whereas ReferFormer and SgMg face difficulties in understanding these complex scenarios. The favorable performance shows that transferring the knowledge of VLP models boosts the vision-language understanding ability of our model.
@@
Figure 13: Qualitative comparisons between Variant-3 (w/o historical prompts) and Variant-4 (w/ historical prompts) on a challenging video. The $55^{th}$ frame is from the previous clip, and both Variant-3 and Variant-4 successfully locate the referred rabbit in this frame. With the historical prior of the referred rabbit, Variant-4 keeps tracking it in the current clip (from the $60^{th}$ to the $85^{th}$ frame). By contrast, the segmentation masks of Variant-3 drift to the distractor in the current clip, which is a similar rabbit gradually appearing in the view.
Figure 14: Prediction results for segmenting a mirror. The referred object is a mirror located on the left above the sink. Although being able to understand the concept of the mirror and locate it, VLP-RVOS has difficulty distinguishing between the reflection inside the mirror and the real object outside, as highlighted by the red circles.
Comparison between w/o and w/ temporal prompts. We qualitatively compare Variant-2 (w/o temporal prompts) and Variant-3 (w/ temporal prompts) to obtain more insights into the effect of the temporal prompt. Figure 12 illustrates their segmentation results over consecutive frames of the same clip on several videos. Note that the videos in Ref-Youtube-VOS are annotated every 5 frames and VLP-RVOS performs segmentation on the annotated frames. We can observe that Variant-3 generates much more stable and consistent segmentation masks across consecutive frames than Variant-2. These comparisons demonstrate the effectiveness of temporal prompts for temporal modeling.
Comparison between w/o and w/ historical prompts. We also qualitatively compare Variant-3 (w/o historical prompts) and Variant-4 (w/ historical prompts) to analyze the effect of the historical prompt. Figure 13 illustrates their segmentation results over two consecutive clips on a video. Both Variant-3 and Variant-4 can locate the referred rabbit in the previous clip (as shown in the $55^{th}$ frame). Nevertheless, Variant-3 loses the referred rabbit and drifts to a distractor in the following clip (from the $60^{th}$ to the $85^{th}$ frame). By contrast, Variant-4 with the target prior from the previous clip continues tracking this rabbit in the following clip.
Failure cases. Although VLP-RVOS has shown promising vision-language understanding abilities in the above experiments, we observe a challenge in distinguishing between reflections inside mirrors and real objects outside. As shown in Figure 14, VLP-RVOS can recognize the presence of the mirror, but it incorrectly identifies the reflection inside the mirror as the real object. A potential and straightforward solution is to further enhance the contextual modeling ability and meanwhile incorporate the mirror data [74] for training.
@@ V Conclusion
We have presented a VLP-RVOS framework to transfer VLP models to RVOS. It enables learning relation modeling for RVOS from aligned VL space instead of from scratch. Specifically, we propose a temporal-aware prompt-tuning method, which not only adapts pre-trained representations for pixel-level prediction but also empowers the vision encoder to model temporal clues. We further propose a multi-stage VL relation modeling scheme for comprehensive VL understanding. Besides, we design a cube-frame attention mechanism for efficient and effective spatial-temporal reasoning. Extensive experiments on four benchmarks demonstrate the effectiveness and generalization of VLP-RVOS.

### v2

@@ IV-B Ablation studies
TABLE III: Experimental results of cross-using VLP models and applying CLIP to other RVOS frameworks. All models are trained on Ref-Youtube-VOS alone.
Algorithms Pretrained Pretrained Aligned $\mathcal{J}\&\mathcal{F}$ (%)
@@
ReferFormer [7]+CLIP CLIP ViT-B/16 CLIP BERT 51.8 51.1
SgMg [68]+CLIP CLIP ViT-B/16 CLIP BERT 52.7 51.9
@@
TABLE IV: Experimental results with varying inference clip lengths of our VLP-RVOS with ViT-B/16 CLIP on Ref-Youtube-VOS. Please note that videos in Ref-Youtube-VOS are 6 FPS.
Clip length
Time duration
3
0.5s
6
1s
12
2s
18
3s
24
4s
30
5s
36
6s
Var.
$\mathcal{J}\&\mathcal{F}$ (%) 62.7 62.9 62.8 63.0 63.1 62.8 62.8 0.019
FLOPs (G) 73.27 71.91 71.25 71.04 70.94 70.89 70.87
GPU Mem. (MB) 2,529 3,051 3,809 4,645 5,063 5,431 6,589
IV-B1 Analyses on proposed components
We analyze the proposed components through 10 variants, as shown in Table I. The experiments begin with a baseline (Variant-1) consisting of a frozen VLP model, $N$ transformer decoder layers originally used for stage-3 VL relation modeling, and a segmentation head.
Analyses on temporal-aware VL prompt-tuning. We gradually introduce different prompts into the baseline to analyze their effect (Variant-2/3/4). The language and vision prompts enable Variant-2 to adapt pretrained representations for pixel-level prediction, improving $\mathcal{J}\&\mathcal{F}$ by 3.9%. By introducing the temporal prompts (3,072 learnable parameters) and the PRTC module (no learnable parameters), Variant-3 improves $\mathcal{J}\&\mathcal{F}$ by 2.3%. It shows that a few learnable parameters effectively enhance the temporal modeling ability. The performance gap between Variant-3 and Variant-4 indicates that historical prompts benefit RVOS in the clip-by-clip inference paradigm.
@@ IV-B1 Analyses on proposed components
Analyses on spatial-temporal attention. We construct Variant-8/9/10 modeling the dense spatial-temporal relation with our CF-MSA, global MSA, and 3D SW/W-MSA [9], respectively. Compared with Variant-7 which only models the temporal context by the vision encoder, all the attention mechanisms bring performance gains, demonstrating the necessity of explicit spatial-temporal relation modeling. Besides, CF-MSA achieves the largest performance gain of 2.2% among the three attention methods, demonstrating its effectiveness.
IV-B2 Analyses on model tuning methods
@@
We conduct experiments with several popular model tuning methods on VLP-RVOS to analyze their effect. The involved methods include: 1) Frozen, in which the VLP model is frozen; 2) Partial-$m$, in which the last $m$ layers of the vision encoder are fine-tuned; 3) Full fine-tuning, in which the entire vision encoder is fine-tuned (note that the language encoder is kept frozen following [7, 28, 15]); 4) Adapter-tuning, in which additional MLP layers are introduced to tune pretrained representations. Herein we use UniAdapter [69], which has proven to be effective on several cross-modality tasks.
@@
We report the within-dataset (Ref-Youtube-VOS) and cross-dataset (Ref-DAVIS17) evaluation results in Table II. Although the frozen method obtains the worst performance on Ref-Youtube-VOS, it performs well on Ref-DAVIS17 as it retains the pretrained knowledge. Compared with partial fine-tuning, full fine-tuning cannot obtain better performance and even harm generalization. UniAdapter [69] obtains mediocre performance. We speculate the reason is that it is designed for image-level VL understanding. Our prompt-tuning performs best on the two benchmarks, demonstrating its effectiveness and generalization.
TABLE V: Experimental results on Ref-YouTube-VOS and Ref-DAVIS17. Many RVOS methods use Video-Swin as the visual backbone, while VLP models typically use ViT as the visual encoder. For relatively fair evaluation, we measure the FLOPs per frame and speed of the RVOS models on RTX3090 and split those with similar efficiency to the same group for comparison. Specifically, we compare our models using ViT-B and VLMo-B with those using Video-Swin-T, and compare our models using ViT-L and VLMo-L with those using Video-Swin-B and Swin-L, considering their similar FLOPs and Speed.
Algorithms Visual Backbone FLOPs (G) Speed (FPS) Ref-YouTube-VOS Ref-DAVIS17
Trained on Ref-Youtube-VOS alone
MTTR [6] Video-Swin-T 55.3 54.0 56.6
MANet [21] Video-Swin-T 55.6 54.8 56.5
ReferFormer [7] Video-Swin-T 72 59 56.0 54.8 57.3 55.8 51.8 59.8
SgMg [68] Video-Swin-T 65 67 58.9 57.7 60.0 56.7 53.3 60.0
SOC [70] Video-Swin-T 43 64 59.2 57.8 60.5 59.0 55.4 62.6
Ours (CLIP) ViT-B/16 72 77 59.7 57.9 61.5 60.3 56.7 64.0
Ours (VLMo) VLMo-B 61 55 60.1 58.4 61.8 61.2 57.3 65.1
Pretrained on Ref-COCO/+/g and fine-tuned on Ref-Youtube-VOS
R2VOS [3] Video-Swin-T 61.3 59.6 63.1
ReferFormer [7] Video-Swin-T 72 59 59.4 58.0 60.9 59.7 56.6 62.8
SgMg [68] Video-Swin-T 65 67 62.0 60.4 63.5 61.9 59.0 64.8
SOC [70] Video-Swin-T 43 64 62.4 61.1 63.7 63.5 60.2 66.7
Ours (CLIP) ViT-B/16 72 77 62.9 61.3 64.4 65.1 61.4 68.8
Ours (VLMo) VLMo-B 61 55 63.1 61.5 64.7 65.5 60.7 69.4
ReferFormer [7] Video-Swin-B 132 35 62.9 61.3 64.6 61.1 58.1 64.1
OnlineRefer [28] Video-Swin-B 127 11 62.9 61.0 64.7 62.4 59.1 65.6
VLT [71] Video-Swin-B 63.8 61.9 65.6 61.6 58.9 64.3
HTML [31] Video-Swin-B 63.4 61.5 65.2 62.1 59.2 65.1
SgMg [68] Video-Swin-B 121 41 65.7 63.9 67.4 63.3 60.6 66.0
TempCD [32] Video-Swin-B 65.8 63.6 68.0 64.6 61.6 67.6
SOC [70] Video-Swin-B 98 34 66.0 64.1 67.9 64.2 61.0 67.4
DsHmp [72] Video-Swin-B 67.1 65.0 69.1 64.9 61.7 68.1
ReferFormer [7] Swin-L 220 37 62.4 60.8 64.0 60.5 57.6 63.4
HTML [31] Swin-L 63.4 61.5 65.3 61.6 58.9 64.4
OnlineRefer [28] Swin-L 222 11 63.5 61.6 65.5 64.8 61.6 67.7
HTR [73] Swin-L 67.1 65.3 68.9 65.6 62.3 68.8
Ours (CLIP) ViT-L/14 219 30 66.0 63.6 68.3 68.2 64.6 71.8
Ours (VLMo) VLMo-L 183 22 67.6 65.3 69.8 70.2 66.3 74.1
IV-B3 Analyses on temporal modeling methods
@@
We conduct experiments with two popular methods also designed for temporal modeling within vision transformers, TeViT [66] and IFC [67]. TeViT [66] shifts several learnable tokens across frames for temporal modeling. IFC [67] introduces a trainable transformer encoder for temporal aggregation, introducing 66 M learnable parameters. We evaluate the two methods in VLP-RVOS and report the results in Table II. The comparisons demonstrate the superiority of our PRTC. Moreover, PRTC occupies a small proportion of the computational load during the visual encoding process. For instance, the total FLOPs per frame for visual encoding with VLMo-L amount to approximately 152.2G, whereas for PRTC, it is only 1.9G.
IV-B4 Effect of the prompt token number
@@
IV-B5 Effect of the aligned VL space and our transferring framework
@@
We delve deeper into analyses by breaking the aligned VL space and applying CLIP to existing RVOS frameworks. We break the aligned VL space by cross-using the vision and language encoders of CLIP and VLMo, which results in significant performance drops, as shown in Table III. We also integrate ViT-B/16 CLIP into ReferFormer and SgMg, where we follow [74] to obtain hierarchical features based on ViT. Herein we use the same input image size as our VLP-RVOS, which is 352. Table III shows that ReferFormer and SgMg using CLIP obtain inferior performance compared with our VLP-RVOS. This implies that both ReferFormer and SgMg fail to fully harness the potential of the pretrained CLIP model for RVOS. Overall, these results highlight that both the aligned VL space and our transferring framework are crucial for VLP-RVOS to achieve state-of-the-art performance.
IV-B6 Temporal modeling across different time spans
We measure $\mathcal{J}\&\mathcal{F}$, FLOPs per frame, and GPU memory usage of our VLP-RVOS with different inference clip lengths. As shown in Table IV, our VLP-RVOS exhibits stable performance with varying inference clip lengths (the variance is 0.019), validating its robustness to clip length and strong temporal modeling ability. We also observe that GPU memory usage increases but the FLOPs decrease as the clip length increases. Users can adjust the inference clip length based on the available hardware memory without worrying about performance degradation in real-world applications.
TABLE VI: Experimental results on A2D/JHMDB-Sentences. All the models are trained on the A2D-Sentences training set alone.
Algorithms Visual Backbone A2D-Sentences JHMDB-Sentences
mAP (%) IoUOverall (%) IoUMean (%) mAP (%) IoUOverall (%) IoUMean (%)
LBDT-4 [4] ResNet-50 47.2 70.4 62.1 41.1 64.5 65.8
TempCD [32] ResNet-50 0. 76.6 68.6 70.6 69.6
LoSh-R [75] Video-Swin-T 50.4 74.3 66.6 40.7 71.6 71.3
SOC [70] Video-Swin-T 50.4 74.7 66.9 39.7 70.7 70.1
ReferFormer [7] Video-Swin-S 53.9 77.7 69.8 42.4 72.8 71.5
OnlineRefer [28] Video-Swin-B 0. 79.6 70.5 73.5 71.9
Ours (CLIP) ViT-B/16 53.3 76.7 69.5 44.2 73.6 71.9
Ours (VLMo) VLMo-B 53.9 78.5 72.7 44.6 73.7 72.3
Ours (CLIP) ViT-L/14 59.4 84.0 75.3 46.0 77.9 75.9
Ours (VLMo) VLMo-L 63.1 86.2 77.7 47.1 78.3 76.6
IV-C Comparison with state-of-the-art methods
Ref-Youtube-VOS & Ref-DAVIS17. Many RVOS methods use Video-Swin as the visual backbone, while ours use ViT as the visual encoder. For relatively fair evaluation, we measure the FLOPs per frame and speed of the RVOS models on RTX3090 and split those with similar efficiency to the same group for comparison. Specifically, we compare our models using ViT-B and VLMo-B with those using Video-Swin-T, and compare our models using ViT-L and VLMo-L with those using Video-Swin-B and Swin-L. Table V reports the results using different training protocols.
On Ref-Youtube-VOS, our models with VLMo achieve the best performance in all metrics in both groups, and our models with CLIP perform comparably with state-of-the-art RVOS methods in the two groups. These comparisons demonstrate the effectiveness of our VLP-RVOS framework. Besides, our models with VLMo and CLIP exhibit substantial advantages on Ref-DAVIS17 compared with other RVOS algorithms. The cross-dataset evaluation, i.e., training on Ref-Youtube-VOS and testing on Ref-DAVIS17, highlights the strong generalization ability of our VLP-RVOS. In terms of efficiency, our models achieve real-time or nearly real-time speeds.
A2D-Sentences & JHMDB-Sentences. Table VI presents the experimental results on A2D-Sentences and JHMDB-Sentences. All the models are trained on A2D-Sentences. Our models using ViT-B/14 CLIP and VLMo-B perform favorably against recently proposed methods using Video-Swin-T/S, such as Losh-R [75], SOC [70] and ReferFormer [7]. Besides, our models using ViT-L/14 CLIP and VLMo-L outperform OnlineRefer [28] using Video-Swin-B by large margins in IoU.
TABLE VII: Experimental results on the MeViS validation set. All the models are trained on the MeViS training set alone.
Algorithms $\mathcal{J}\&\mathcal{F}$ (%) $\mathcal{J}$ (%) $\mathcal{F}$ (%)
URVOS [18] 27.8 25.7 29.9
LBDT [4] 29.3 27.8 30.8
MTTR [6] 30.0 28.8 31.2
ReferFormer [7] 31.0 29.8 32.2
VLT+TC [71] 35.5 33.6 37.3
LMPM [19] 37.2 34.2 40.2
DsHmp [72] 46.4 43.0 49.8
Ours (ViT-B/16 CLIP) 44.6 41.3 48.0
Ours (VLMo-B) 45.4 42.0 48.8
MeViS. MeViS [19] is a benchmark requiring RVOS models to understand the motion in video to locate and segment the target object. We conduct experiments on MeViS to evaluate the motion modeling ability of VLP-RVOS. Table VII reports the results. LMPM [19] and DsHmp [72] are two RVOS algorithms elaborated to comprehend the motion of the target object with a motion perception mechanism, which achieve astonishing progress on MeViS. Although without explicit motion modeling at the object level, our models obtain comparable performance with DsHmp and better performance than LMPM, demonstrating its effectiveness in modeling the temporal context within the video clip.
@@ IV-D Qualitative results
Figure 10 shows the segmentation results in a video with cluttered scenes. Our VLP-RVOS accurately comprehends detailed descriptions and precisely segments the target objects, whereas ReferFormer and SgMg face difficulties in understanding these complex scenarios. The favorable performance shows that transferring the knowledge of VLP models boosts the vision-language understanding ability of our model.
@@
Figure 12: Qualitative comparisons between Variant-3 (w/o historical prompts) and Variant-4 (w/ historical prompts) on a challenging video. The $55^{th}$ frame is from the previous clip, and both Variant-3 and Variant-4 successfully locate the target rabbit in this frame. With the historical prior of the target rabbit, Variant-4 keeps tracking it in the current clip (from the $60^{th}$ to the $85^{th}$ frame). By contrast, the segmentation masks of Variant-3 drift to the distractor in the current clip, which is a similar rabbit gradually appearing in the view.
Figure 13: Prediction results for segmenting a mirror. The target object is a mirror located on the left above the sink. Although being able to understand the concept of the mirror and locate it, VLP-RVOS has difficulty distinguishing between the reflection inside the mirror and the real object outside, as highlighted by the red circles.
Comparison between w/o and w/ temporal prompts. We qualitatively compare Variant-2 (w/o temporal prompts) and Variant-3 (w/ temporal prompts) to obtain more insights into the effect of the temporal prompt. Figure 11 illustrates their segmentation results over consecutive frames of the same clip on several videos. Note that the videos in Ref-Youtube-VOS are annotated every 5 frames and VLP-RVOS performs segmentation on the annotated frames. We can observe that Variant-3 generates much more stable and consistent segmentation masks across consecutive frames than Variant-2. These comparisons demonstrate the effectiveness of the temporal prompts for temporal modeling.
Comparison between w/o and w/ historical prompts. We also qualitatively compare Variant-3 (w/o historical prompts) and Variant-4 (w/ historical prompts) to analyze the effect of the historical prompt. Figure 12 illustrates their segmentation results over two consecutive clips on a video. Both Variant-3 and Variant-4 can locate the target rabbit in the previous clip (as shown in the $55^{th}$ frame). Nevertheless, Variant-3 loses the referred rabbit and drifts to a distractor in the following clip (from the $60^{th}$ to the $85^{th}$ frame). By contrast, Variant-4 with the target prior from the previous clip continues tracking this rabbit in the following clip.
Failure cases. Although VLP-RVOS has shown promising vision-language understanding abilities in the above experiments, we observe a challenge in distinguishing between reflections inside mirrors and real objects outside. As shown in Figure 13, VLP-RVOS can recognize the presence of the target mirror, but it incorrectly identifies the reflection inside the mirror as the real object. A potential and straightforward solution is to further enhance the contextual modeling ability and meanwhile incorporate the mirror data [76] for training.
@@ V Conclusion
We have presented a VLP-RVOS framework to transfer VLP models to RVOS. It enables learning relation modeling for RVOS from aligned VL space instead of from scratch. Specifically, we propose a temporal-aware prompt-tuning method, which not only adapts pre-trained representations for pixel-level prediction but also empowers the vision encoder to model temporal clues. We further design a cube-frame attention mechanism for efficient and effective spatial-temporal reasoning. Besides, we propose a multi-stage VL relation modeling scheme for comprehensive VL understanding. Extensive experiments on four benchmarks demonstrate the effectiveness and generalization of VLP-RVOS.